OpenStudy (anonymous):

Solve the system using the substitution method x-10y=30 3y-4x=-46

sam (.sam.):

x-10y=30 x=30+10y ---------------------- 3y-4(30+10y)=-46 3y-120-40y=-46 -37y-120=-46 -37y=74 y=-2 ---------------------- x-10(-2)=30 x+20=30 x=10